OpenStudy (anonymous):

Solve. 1 = –(–w + 6) A. –7 B. –3 C. 7 D. 3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

so the two - sighs make a + right?

OpenStudy (skullpatrol):

Yes, $$(-1)(-1)=1$$

OpenStudy (anonymous):

then I have no clue O_O

OpenStudy (skullpatrol):

recall $$-a = (-1)a$$

OpenStudy (anonymous):

okay.. well You just confused me further..

OpenStudy (skullpatrol):

consider $$a=-w+6$$

OpenStudy (skullpatrol):

$$-(-w+6)=-1(-w+6)$$

OpenStudy (skullpatrol):

$$-1(-w+6)=(-1)(-w)+(-1)(6)=w-6$$

OpenStudy (skullpatrol):

To find the opposite of a sum or difference, you change the sign of each term of the sum or difference.

OpenStudy (skullpatrol):

$$-(-w+6)=w-6$$

OpenStudy (anonymous):

is it c? :/

OpenStudy (skullpatrol):

Yes.